The common stock of NCP paid ​$1.45 in dividends last year. Dividends are expected to grow at an annual rate of 6.40 percent for an indefinite number of years.

a. If​ NCP's current market price is ​$23.05 per​ share, what is the​ stock's expected rate of​ return?

b. If your required rate of return is 8.4 ​percent, what is the value of the stock for​ you?

c. Should you make the​ investment?
Answer #1

A. Expected rate of return [k] = D1 / P0 +g%

=1.45*1.0640 / 23.05 +6.4%

=1.5428/23.05 +6.4%

=13.09%

B. Value of stock = D1 / K -g

=1.45*1.0640 / 8.4-6.4

=1.5428 / 2%

=77.14

C.Yes,reason being value of stock is higher than current market value.